OpenStudy (wolfcherry):

How would you describe the interaction between opposite (N-S or S-N) magnetic poles? They repel each other. They are attracted to each other. They create a magnetic field from S to N. There is repulsion between N and S poles. No magnetic field exists between the magnets.

OpenStudy (isaidavila):

They will be attracted to each others

OpenStudy (shamim):

U know Like poles repel

OpenStudy (shamim):

Unlike poles attract

OpenStudy (shamim):

N-N=like poles

OpenStudy (shamim):

S-S=like poles

OpenStudy (shamim):

N-S=unlike poles

OpenStudy (shamim):

S-N=unlike poles